solve using the addition method
x+y=10
x-y=8

Answer by nerdybill(7384)

Starting with:
x+y=10
x-y=8
.
We simply ADD both equations together:
x+y=10
x-y=8
---------
2x = 18
x = 9
.
substitute the value of x above into either of the two original equations to find y:
x+y=10
9+y=10
y = 1
.
Solution: (x,y) = (9,1)